Circular No. 7161

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V1333 AQUILAE R. Jain and C. Bailyn, Yale University; M. Garcia and K. Rines,

Harvard-Smithsonian Center for Astrophysics; A. Levine, Massachusetts Institute of Technology; and J. Espinoza and D. Gonzalez, Cerro Tololo Interamerican Observatory (CTIO), report YALO consortium observations: “Aquila X-1 = V1333 Aql appears to be beginning a new outburst. This x-ray binary outbursts approximately once per year, and based on its recent outbursts was due to erupt. We have monitored the optical counterpart of the soft x-ray transient Aquila X-1 since early Mar. in Johnson R, using the YALO 1-m telescope at CTIO. Compared to the average quiescent level, which has a fluctuation of < 0.1 mag, the source had brightened in R by ∼ 0.2 mag on May 8.3 UT. During May 10.2–10.38, we obtained five data points spanning 9 hr and found that the brightness increased from 0.4 to 0.55 mag brighter than the quiescent average. V -band data from the 1.22m telescope at the Whipple Observatory show similar increases. We find that by May 9.5 the source had brightened by ∼ 0.4 mag in V, compared to the average quiescent level. Subsequently we obtained three YALO V images on May 10.2–10.38 and found an increase in brightness of ∼ 0.45 to 0.6 mag, again compared to the quiescent average. We will continue monitoring this source both at optical and at x-ray (using RXTE) wavelengths. Observations at all wavelengths are encouraged.”

SUPERNOVA 1999ca IN NGC 3120 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Center for Astro-

physics, report that a spectrum of SN 1999ca was obtained by M. Calkins with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on May 7.2 UT, showing it to be a type-II supernova after maximum light. The spectrum exhibits a P-Cyg Hα profile and absorption features from Hβ and Fe II, and is similar to a spectrum of the type-II SN 1979C several weeks after maximum light (Branch et al. 1981, Ap.J. 244, 780). Adopting the NASA/IPAC Extragalactic Database recession velocity for NGC 3120 of 2791 km/s implies a photospheric expansion velocity of 7900 km/s for SN 1999ca, measured using the Balmer-line absorption troughs.

COMET C/1999 H1 (LEE) Visual m_{1} estimates: May 5.44 UT, 7.7 (C. E. Drescher, Warrill View,

Queensland, 10×50 binoculars); 9.16, 7.0 (R. Keen, Mt. Thorodin, CO, 7×50 binoc.); 10.33, 7.2 (S. T. Rae, Kopane, New Zealand, 10×50 binoc.).

1999 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green